OpenStudy (jkbo):

Please help Determine whether the function is linear or quadratic. Identify the quadratic, linear, and constant terms. f(x)= (3x-4)(-6x+5) A.quadratic function quadratic term: -18x^2 linear term: 39x constant term: –20 B.linear function linear term: -18x^2 constant term: –20 C.quadratic function quadratic term: -12x^2 linear term: -42x constant term: –20 D.linear function linear term: 39x constant term: –20

OpenStudy (anonymous):

Do you know how to distribute multiplication? e.g. \((a+b)(c+d)=a(c+d)+b(c+d)=ac+ad+bc+bd\)

OpenStudy (jkbo):

yes I figured out he answer